# Model Stock: All we need is just a few fine-tuned models

### v1

@@ 6 Conclusion
  • We present experimental settings of §4 and additional experiments of Model Stock in §0.E.
@@ Appendix 0.A Angle and Norm Consistency
We argue that, as discussed in §2.1, angles and norms of fine-tuned weights would remain consistent across fine-tuned models, independent of various factors. These factors include architecture type (ViTs [5], ResNet [9], ConvNeXt [21]), optimizers (SGD, AdamW [22]), augmentations (RRC [27], RandAug [4]), datasets (CIFAR [14], ImageNet [26]), or the initialization of the classifier (zero-shot, LP as in LP-FT [15]). We depict the layer-wise angle and norm of 5 fine-tuned weights for each category based on different random seeds. We give detailed illustrations for each setting at the end of the Appendix to enhance readability (refer to Fig. FM). Across all these settings, the angle and norm of weights exhibit a surprising level of consistency.
@@ 0.A.1 Analysis on layer-wise tendency
The layer-wise angle and norm across various settings are shown in Fig. FJ. We visualize with every weight of attentions/convolutions (Attention/Conv), multi-layer perceptrons (MLP), normalizations (LayerNorm and BatchNorm), a classifier (Classifier), individual bias (Bias), and the remaining layers (i.e., the patchification layer, positional embedding, class embedding, and projection layer). We further display All in each figure, which denotes the concatenation of the weights of entire layers.
@@ 0.A.2 Maintaining consistency during training
We further argue that the consistency we observed is maintained while training progresses, as illustrated by multiple thin shells in Fig. 5. To demonstrate that the angle and norm of fine-tuned models remain consistent during the entire training process, we plot their relationship across weights for every epoch in Fig. K. Please note that the angle is consistent across differently seeded models at the same timestamp (i.e., $\mathbf{w}_{1}|_{t=t_{1}}$ and $\mathbf{w}_{2}|_{t=t_{1}}$), not across models at different timestamps (i.e., $\mathbf{w}_{1}|_{t=t_{1}}$ and $\mathbf{w}_{1}|_{t=t_{2}}$). The observed trend is as follows: as training progresses, the angle between weights steadily decreases. This analysis uses the CLIP ViT-B/32 model fine-tuned on ImageNet-1K with five random seeds.
@@ 0.A.3 Filter-wise analysis of weights
Li et al. [18] showed that when evaluating the robustness of a neural network by adding random noise to certain weights, performance analysis based on adding filter-wise noise (i.e., adding noise for each row in all weight matrices) aligns more closely with the generalization performance than adding layer-wise noise does. Inspired by this observation, we investigate the possibility that the weight distribution may follow a filter-wise Gaussian distribution and adapt this concept to our method (see the performance analysis in §0.E.4). Fig. L illustrates the angle distribution filter-wise. The angle exhibits much larger standard deviations than the layer-wise distribution. This could be attributed to the reduction in dimensionality. As the number of dimensions decreases, it becomes challenging to approximate the norm as a constant value.
@@ 0.A.4 Analysis on non-CLIP models
To verify if this key observation also applies to non-CLIP models, we analyze the geometric patterns of fine-tuned weights trained using the DeiT [30] method (i.e., pre-trained on ImageNet-21K). Fig. M displays the angle and norm of 10 DeiT-base models first pre-trained on ImageNet-21K [26] and then fine-tuned on ImageNet-1K. We find that weights pre-trained with ImageNet-21K also exhibit consistent angle and norm, indicating that our observation may be valid beyond CLIP fine-tuning scenarios as well.
@@ Appendix 0.D Detailed Proof of Model Stock
Appendix 0.E Additional Experiments
Here, we present detailed setups for the experiments in §4 and additional experimental studies to verify the effectiveness and applicability of Model Stock.
0.E.1 Experimental setup
@@
We utilize AdamW optimizer [22] with a weight decay of 0.1. We employ two training setups for Model Stock. The first is training Model Stock with a learning rate of $3\times 10^{-5}$ in 10 epochs with minimal data augmentation. The minimal data augmentation utilizes random resize crop augmentation with a minimum crop ratio of $0.9$, mixup [35] augmentation with $\beta$=0.5, following Model Soup’s “standard grid search” setting. The other is training Model Stock with a learning rate of $2\times 10^{-5}$ in 16 epochs with strong data augmentation. The strong data augmentation utilizes random resize crop augmentation with a minimum crop ratio of $0.08$ and random augmentation [4] ($N=2$, $M=10$) following Model Soup’s “random search” setting. When experimenting with the ViT-B/16 and ViT-L/14 models, we adjusted the learning rate and batch size to accommodate the GPU memory constraints.
@@
Figure E: Results on LP initialization. We plot in-distribution ImageNet accuracy (x-axis) and distribution shift results (y-axis) with individual fine-tuned models (gray circles) and Model Soups [32]. Note that Model Stock has much smaller ($35\times$ smaller) computational costs than Model Soups, leveraging 71 various fine-tuned models as in the original paper.
@@
Table A: Complete results of Table 3 with ObjectNet [1] and ImageNet-ReaL [2].
@@
0.E.2 Experiments with LP initialization
We conduct Model Stock with LP initialization and compare it with Model Soups that are initialized from LP. The results are in Fig. E. In this experiment, we use the 16-epoch training setup with strong data augmentation for training Model Stock. As shown in Fig. E, Model Stock outperforms the individual fine-tuned models [^6] All the individual model checkpoints are from the official Model Soup repository. (gray dots) on ImageNet accuracy. Model Stock also demonstrates competitive performance against Model Soups considering WiSE-FT curves. Note that Model Stock is much more efficient ($35\times$) than Model Soups, which utilize $71$ models in this experiment.
0.E.3 Complete results with ObjectNet benchmark
In the main paper, we omit the results of ObjectNet [1] on CLIP ViT-B/16 experiments since the comparison methods such as LP-FT [15], FTP [29] have not evaluated on ObjectNet benchmark. We here show the results with ObjectNet [1] and ImageNet-ReaL [2] of CLIP ViT-B/16 in Table A. Model Stock consistently demonstrates its effectiveness with ObjectNet and ImageNet-ReaL as well.
Table B: Performance comparison of merging units in Model Stock. This table presents the overall performance of Model Stock using different merging units: entire weight merging, entire weight merging based on transformer block angle, layer-wise merging, and filter-wise merging. It highlights the effectiveness of each strategy in approaching the weight center and their impact on the model’s performance.
Merging Unit Target Dist. Shifts
@@ 0.E.3 Complete results with ObjectNet benchmark
0.E.4 Ablation study on merging unit
We investigate the efficacy of different merging units within our method, Model Stock. Our default approach employs layer-wise merging, but alternatives include merging based on the angle between 1) entire weights, 2) weights of the entire repetitive transformer blocks following [32], or 3) using a filter-wise approach as discussed in §0.A.3. The results of these ablations are summarized in Table B, where we assess the overall performance based on the chosen merging unit.
Our analysis reveals that the accuracy of noise distribution estimation is critical in approaching the weight center. When assuming weight noise across the entire model, our method does not approximate the weight center as effectively as it does with layer-wise merging, leading to suboptimal overall performance. Similarly, the merging performance based on the angle of transformer blocks was insufficient. Conversely, while filter-wise noise demonstrates a larger standard deviation in angle, as depicted in Fig. L, this increased variance results in a more significant error in Gaussian distribution approximation. Consequently, the overall performance under filter-wise merging is slightly inferior to layer-wise one.
@@ 0.E.4 Ablation study on merging unit
Figure F: Layer-wise angle and norm across different model architectures. The angle and norm for CLIP ViT-L/14, CLIP ResNet50, and OpenCLIP ConvNeXt are displayed from top to bottom. These metrics demonstrate consistency regardless of the model type from left (first layer) to right (last layer). It is important to note that we also depict the error bars for each layer in all figures, but they are not visible in most layers due to the small standard deviation.
@@
Figure G: Layer-wise angle and norm across different optimizers. Displayed from top to bottom are the angle and norm for models trained with SGD and SGD with momentum, respectively. These metrics demonstrate consistency regardless of the optimization strategy from left (first layer) to right (last layer).
@@
Figure H: Layer-wise angle and norm across different augmentations. Displayed from top to bottom are the angle and norm for the vanilla model (10 epochs + no augmentation), +longer epochs (16 epochs), and +RRC. Each augmentation is applied incrementally. These metrics demonstrate consistency regardless of the augmentations from left (first layer) to right (last layer).
Figure I: Layer-wise angle and norm across different datasets. The angle and norm for models trained on different datasets, including CIFAR [14] are displayed from top to bottom. These metrics demonstrate consistency regardless of the dataset type from left (first layer) to right (last layer).
Figure J: Layer-wise angle and norm across different classifier initializations. The angle and norm for models trained with differently initialized networks following the LP-FT [15] method are displayed from top to bottom. These metrics demonstrate consistency regardless of the initialization method from left (first layer) to right (last layer).
Figure K: Layer-wise angle during training. Displayed are the overlapped angles across models trained with different random seeds at each timestamp. Even during training, the angle remains highly consistent, decreasing as training progresses.
@@
Figure L: Filter-wise angle for attention and MLP layers in ViT-B/32. We display filter-wise angles for each layer. Each bar represents each row (i.e., filter) in the given layer. Interestingly, the angles between the filters of the fine-tuned weights exhibit similar values, while the standard deviation between each filter is notably larger than that of the angle between each layer. Due to the large number of layers, only representative layers are selected for display.
Figure M: Layer-wise angle and norm for DeiT. The angle and norm for DeiT-base models are displayed, each trained with different random seeds. These models are initially pre-trained on ImageNet-21K [26] and then fine-tuned on ImageNet-1K. The consistency observed in the metrics is maintained even in the DeiT training setting.

### v2

@@ 7 Acknowledgment
  • We discuss prior studies through the lens of our findings in §0.E.
  • We provide an additional analysis of the interpolation ratio in §0.F.
  • We present experimental settings of §4 in §0.G.
  • We present additional experiments of Model Stock in §0.H
@@ Appendix 0.A Angle and Norm Consistency
We argue that, as discussed in §2.1, angles and norms of fine-tuned weights would remain consistent across fine-tuned models, independent of various factors. These factors include architecture type (ViTs [5], ResNet [9], ConvNeXt [24]), optimizers (SGD, AdamW [25]), augmentations (RRC [35], RandAug [4]), datasets (CIFAR [17], ImageNet [34]), or the initialization of the classifier (zero-shot, LP as in LP-FT [18]). We depict the layer-wise angle and norm of 5 fine-tuned weights for each category based on different random seeds. We give detailed illustrations for each setting at the end of the Appendix to enhance readability (refer to Fig. HO). Across all these settings, the angle and norm of weights exhibit a surprising level of consistency.
@@ 0.A.1 Analysis on layer-wise tendency
The layer-wise angle and norm across various settings are shown in Fig. HL. We visualize with every weight of attentions/convolutions (Attention/Conv), multi-layer perceptrons (MLP), normalizations (LayerNorm and BatchNorm), a classifier (Classifier), individual bias (Bias), and the remaining layers (i.e., the patchification layer, positional embedding, class embedding, and projection layer). We further display All in each figure, which denotes the concatenation of the weights of entire layers.
@@ 0.A.2 Maintaining consistency during training
We further argue that the consistency we observed is maintained while training progresses, as illustrated by multiple thin shells in Fig. 5. To demonstrate that the angle and norm of fine-tuned models remain consistent during the entire training process, we plot their relationship across weights for every epoch in Fig. M. Please note that the angle is consistent across differently seeded models at the same timestamp (i.e., $\mathbf{w}_{1}|_{t=t_{1}}$ and $\mathbf{w}_{2}|_{t=t_{1}}$), not across models at different timestamps (i.e., $\mathbf{w}_{1}|_{t=t_{1}}$ and $\mathbf{w}_{1}|_{t=t_{2}}$). The observed trend is as follows: as training progresses, the angle between weights steadily decreases. This analysis uses the CLIP ViT-B/32 model fine-tuned on ImageNet-1K with five random seeds.
@@ 0.A.3 Filter-wise analysis of weights
Li et al. [21] showed that when evaluating the robustness of a neural network by adding random noise to certain weights, performance analysis based on adding filter-wise noise (i.e., adding noise for each row in all weight matrices) aligns more closely with the generalization performance than adding layer-wise noise does. Inspired by this observation, we investigate the possibility that the weight distribution may follow a filter-wise Gaussian distribution and adapt this concept to our method (see the performance analysis in §0.H.5). Fig. N illustrates the angle distribution filter-wise. The angle exhibits much larger standard deviations than the layer-wise distribution. This could be attributed to the reduction in dimensionality. As the number of dimensions decreases, it becomes challenging to approximate the norm as a constant value.
@@ 0.A.4 Analysis on non-CLIP models
To verify if this key observation also applies to non-CLIP models, we analyze the geometric patterns of fine-tuned weights trained using the DeiT [38] method (i.e., pre-trained on ImageNet-21K). Fig. O displays the angle and norm of 10 DeiT-base models first pre-trained on ImageNet-21K [34] and then fine-tuned on ImageNet-1K. We find that weights pre-trained with ImageNet-21K also exhibit consistent angle and norm, indicating that our observation may be valid beyond CLIP fine-tuning scenarios as well.
@@ Appendix 0.D Detailed Proof of Model Stock
Appendix 0.E Discussion — Rethinking Pivotal Prior Studies
Figure E: Ensembling impact disappears when interpolating between two averaged weights. We plot the ImageNet performance of interpolated weights between two selected fine-tuned models in Model Soup [40] (left) and between their corresponding weight centers (right).
In this section, we extend our findings to reinterpret the underlying mechanics in prior studies, WiSE-FT [41] and Model Soups [40], through a consistent rationale to illuminate their effectiveness.
0.E.0.1 WiSE-FT
[41] is a state-of-the-art robust fine-tuning method for CLIP-based models. It demonstrates that linearly combining weights of the pre-trained and fine-tuned models achieves a significant accuracy gain on distribution shifts. We argue that the WiSE-FT model’s superiority over a fine-tuned model can be interpreted by its weights being closer to the center of the corresponding weight distribution. Fig. 3 already showed fine-tuned models typically lie on the periphery of flat minima. Given that the angle $\angle\textbf{w}_{0}\textbf{w}_{\text{avr}}^{(50)}\textbf{w}_{1}$ is nearly a right angle, along the line $\overline{\textbf{w}_{0}\textbf{w}_{1}}$, multiple weight points are closer to the center than a single fine-tuned model, thereby enhancing performance. Note that $\textbf{w}_{H}$ is the closest to the center among the line $\overline{\textbf{w}_{0}\textbf{w}_{1}}$. More discussions on performance boosts observed in distribution shifts are provided in the Appendix 0.C.
0.E.0.2 Model Soup
[40] merges various fine-tuned models’ weights trained from varied hyper-parameters. It has been credited with delivering enhanced performance across ImageNet and distribution shifts. Here, we interpret the performance improvements of Model Soup as the result of the proximity to the center of weight distribution. Consider two weight vectors, $\mathbf{w}_{A}$ and $\mathbf{w}_{B}$, fine-tuned with different hyper-parameters and following Gaussian distribution $\mathcal{N}(\bm{\mu_{A}},\Sigma_{A})$ and $\mathcal{N}(\bm{\mu_{B}},\Sigma_{B})$ respectively. Then, the interpolated weight vector $\mathbf{w}_{AB}=t\cdot\mathbf{w}_{A}+(1-t)\cdot\mathbf{w}_{B}$ also follows a Gaussian distribution $\mathcal{N}(\bm{\mu_{AB}},\Sigma_{AB})$. The expected squared distance from the interpolated weight vector to its mean $\bm{\mu_{AB}}$ is minimized to $\frac{\text{trace}(\Sigma_{A})\text{trace}(\Sigma_{B})}{\text{trace}(\Sigma_{A})+\text{trace}(\Sigma_{B})}$ when $t$ is chosen to $\frac{\text{trace}(\Sigma_{B})}{\text{trace}(\Sigma_{A})+\text{trace}(\Sigma_{B})}$, indicating the reduction of variance through weight interpolation (i.e., the distance between $\mathbf{w}_{AB}$ and $\bm{\mu_{AB}}$ might be closer than each weight’s distance). For example, if $\text{trace}(\Sigma_{B})$ is equal to $\text{trace}(\Sigma_{A})$, this minimum squared distance is exactly half of the sum of the individual traces when $t=0.5$. This insight suggests that the performance gains realized by Model Soup could be due to reduced variance resulting from merging numerous weights.
We set up a toy experiment to evaluate the effect of variance reduction in the Model Soup scenario by comparing the interpolation of fine-tuned weights with the interpolation of their corresponding weight centers, when $N=2$. In the former case, variance reduction exists along with the effect of merging diverse hyper-parameters, while in the latter case, performance gain would only come from hyper-parameter diversity. If the diversity of hyper-parameters is a major factor, the performance gain from interpolation of central weights should remain the same. To test this, we assessed the ImageNet performance of interpolated weights between pairs of fine-tuned models within Greedy Model Soup [^5] We opt for Greedy Model Soup to show that even the interpolation of models from the best merging combination does not benefit from the impact of weight diversity. [40] and compared it to interpolations between their central weights, calculated as the average of 20 differently seeded models. Fig. E shows that, unlike interpolations between individual models, using the centers does not significantly improve performance. This suggests that proximity to the center of the weight distribution may play a more critical role than hyper-parameter diversity in weight ensemble methods in this case.
It is also worth noting that $\bm{\mu_{AB}}$ always surpasses the performance of $\mathbf{w_{AB}}$ for the same interpolation ratio $t$, indicating that the importance of proximity to the center remains consistent for interpolated weights. With extensive future research, this understanding could provide valuable insights for developing more generalizable and effective weight-merging techniques.
(a) Layer depth vs. interpolation ratio $t$.
(b) Epoch vs. interpolation ratio $t$.
Figure F: Trend of interpolation ratio $t$ during Model Stock training.
Appendix 0.F Analysis of the interpolation ratio $t$
We analyze the interpolation ratio $t=\frac{2\cos\theta}{1+\cos\theta}$ in a layer-wise manner. During a Model Stock experiment on CLIP ViT-B/32 with 16 epoch training, we log the layer-wise merge ratios at every merging period. Figure F visualizes the averaged interpolation ratio during Model Stock training. We plot two trends of the interpolation ratio for the layer depth and training step. Our overall observation indicates the bias layers have high merge ratios $t$ ($\simeq 1$) with small angles $\theta$ ($\simeq 0$), implying that the bias layers do not need to enjoy the pre-trained model, similar to our discussion in §2 and §3. Focusing on the weight layers, Figure 6(a) shows a U-shape tendency as the layer depth increases, implying the weights of intermediate layers can be more diverse (i.e., larger angle $\theta$) than those of early and later layers. Our intuition here is that since the early and later layers are directly connected to input data and output labels, respectively, they may not demand the advantage of the pre-trained weight. Figure 6(b) presents that the models at the early training stage are more diverse and they enjoy the pre-trained weights more than those of the later training stage. As the model approaches convergence, the diversity of fine-tuning models decreases (i.e., smaller angle $\theta$).
Appendix 0.G Experimental setup
@@
Here, we present detailed setups for the experiments in §4. We utilize AdamW optimizer [25] with a weight decay of 0.1. We employ two training setups for Model Stock. The first is training Model Stock with a learning rate of $3\times 10^{-5}$ in 10 epochs with minimal data augmentation. The minimal data augmentation utilizes random resize crop augmentation with a minimum crop ratio of $0.9$, mixup [44] augmentation with $\beta$=0.5, following Model Soup’s “standard grid search” setting. The other is training Model Stock with a learning rate of $2\times 10^{-5}$ in 16 epochs with strong data augmentation. The strong data augmentation utilizes random resize crop augmentation with a minimum crop ratio of $0.08$ and random augmentation [4] ($N=2$, $M=10$) following Model Soup’s “random search” setting. When experimenting with the ViT-B/16 and ViT-L/14 models, we adjusted the learning rate and batch size to accommodate the GPU memory constraints.
@@
Figure G: Results on LP initialization. We plot in-distribution ImageNet accuracy (x-axis) and distribution shift results (y-axis) with individual fine-tuned models (gray circles) and Model Soups [40]. Note that Model Stock has much smaller ($35\times$ smaller) computational costs than Model Soups, leveraging 71 various fine-tuned models as in the original paper.
Appendix 0.H Additional Experiments
We present additional experimental studies to verify the effectiveness and applicability of Model Stock.
0.H.1 Experiments with LP initialization
We conduct Model Stock with LP initialization and compare it with Model Soups that are initialized from LP. The results are in Fig. G. In this experiment, we use the 16-epoch training setup with strong data augmentation for training Model Stock. As shown in Fig. G, Model Stock outperforms the individual fine-tuned models [^6] All the individual model checkpoints are from the official Model Soup repository. (gray dots) on ImageNet accuracy. Model Stock also demonstrates competitive performance against Model Soups considering WiSE-FT curves. Note that Model Stock is much more efficient ($35\times$) than Model Soups, which utilize $71$ models in this experiment.
@@ 0.H.1 Experiments with LP initialization
Table A: Complete results of Table LABEL:exp:vit_b_16 with ObjectNet [1] and ImageNet-ReaL [2].
@@
0.H.2 Complete comparison results on CLIP ViT-B/16
In the main paper, we omit the results of ObjectNet [1] on CLIP ViT-B/16 experiments since the comparison methods such as LP-FT [18], FTP [37] have not evaluated on ObjectNet benchmark. We here show the results with ObjectNet [1] and ImageNet-ReaL [2] of CLIP ViT-B/16 in Table A. We additionally compare Model Stock with recent fine-tuning methods including FLYP [7], Lipsum-FT [28], and CaRot [29] Model Stock consistently demonstrates its effectiveness with ObjectNet and ImageNet-ReaL as well.
Table B: Comparison against Model Soups [40] on CLIP ViT-B/16. Model Stock shows comparable performance with Model Soups.
Method ImageNet Avg. shifts
CLIP zero-shot Init. 68.3 58.4
Vanilla FT 82.8 56.6
Vanilla FT 83.7 55.9
Uniform Model Soup 84.4 62.7
Greedy Model Soup 84.3 60.4
Model Stock 84.1 61.0
Model Stock 85.2 58.5
0.H.3 Model Stock vs. Model Soups on CLIP ViT-B/16
[…]
Table C: Model Stock with different hyper-parameters on CLIP ViT-B/32.
Method ImageNet Avg. shifts
Model Stock 79.89 50.99
[…]
0.H.4 Model Stock with different hyper-parameters
To verify the validity of Model Stock beyond the setup of the main paper (i.e., different random seeds with the same hyper-parameters), we conduct Model Stock with different hyper-parameters. In detail, when we fine-tune two models for Model Stock, we choose different hyper-parameter for each model (e.g., learning rate, data augmentation.). To ensure the basic assumption of Model Stock, we use the same batch size and training epochs. C shows the experimental results on CLIP ViT-B/32. We repeat 5 runs and report accuracy with standard deviation. Model Stock with different hyper-parameters shows comparable performance to the original one.
Table D: Performance comparison of merging units in Model Stock. This table presents the overall performance of Model Stock using different merging units: entire weight merging, entire weight merging based on transformer block angle, layer-wise merging, and filter-wise merging. It highlights the effectiveness of each strategy in approaching the weight center and their impact on the model’s performance.
Merging Unit Target Avg. Shifts
@@ 0.H.4 Model Stock with different hyper-parameters
0.H.5 Ablation study on merging unit
We investigate the efficacy of different merging units within our method, Model Stock. Our default approach employs layer-wise merging, but alternatives include merging based on the angle between 1) entire weights, 2) weights of the entire repetitive transformer blocks following [40], or 3) using a filter-wise approach as discussed in §0.A.3. The results of these ablations are summarized in Table D, where we assess the overall performance based on the chosen merging unit.
Our analysis reveals that the accuracy of noise distribution estimation is critical in approaching the weight center. When assuming weight noise across the entire model, our method does not approximate the weight center as effectively as it does with layer-wise merging, leading to suboptimal overall performance. Similarly, the merging performance based on the angle of transformer blocks was insufficient. Conversely, while filter-wise noise demonstrates a larger standard deviation in angle, as depicted in Fig. N, this increased variance results in a more significant error in Gaussian distribution approximation. Consequently, the overall performance under filter-wise merging is slightly inferior to layer-wise one.
@@ 0.H.5 Ablation study on merging unit
Figure H: Layer-wise angle and norm across different model architectures. The angle and norm for CLIP ViT-L/14, CLIP ResNet50, and OpenCLIP ConvNeXt are displayed from top to bottom. These metrics demonstrate consistency regardless of the model type from left (first layer) to right (last layer). It is important to note that we also depict the error bars for each layer in all figures, but they are not visible in most layers due to the small standard deviation.
@@
Figure I: Layer-wise angle and norm across different optimizers. Displayed from top to bottom are the angle and norm for models trained with SGD and SGD with momentum, respectively. These metrics demonstrate consistency regardless of the optimization strategy from left (first layer) to right (last layer).
@@
Figure J: Layer-wise angle and norm across different augmentations. Displayed from top to bottom are the angle and norm for the vanilla model (10 epochs + no augmentation), +longer epochs (16 epochs), and +RRC. Each augmentation is applied incrementally. These metrics demonstrate consistency regardless of the augmentations from left (first layer) to right (last layer).
Figure K: Layer-wise angle and norm across different datasets. The angle and norm for models trained on different datasets, including CIFAR [17] are displayed from top to bottom. These metrics demonstrate consistency regardless of the dataset type from left (first layer) to right (last layer).
Figure L: Layer-wise angle and norm across different classifier initializations. The angle and norm for models trained with differently initialized networks following the LP-FT [18] method are displayed from top to bottom. These metrics demonstrate consistency regardless of the initialization method from left (first layer) to right (last layer).
Figure M: Layer-wise angle during training. Displayed are the overlapped angles across models trained with different random seeds at each timestamp. Even during training, the angle remains highly consistent, decreasing as training progresses.
@@
Figure N: Filter-wise angle for attention and MLP layers in ViT-B/32. We display filter-wise angles for each layer. Each bar represents each row (i.e., filter) in the given layer. Interestingly, the angles between the filters of the fine-tuned weights exhibit similar values, while the standard deviation between each filter is notably larger than that of the angle between each layer. Due to the large number of layers, only representative layers are selected for display.
Figure O: Layer-wise angle and norm for DeiT. The angle and norm for DeiT-base models are displayed, each trained with different random seeds. These models are initially pre-trained on ImageNet-21K [34] and then fine-tuned on ImageNet-1K. The consistency observed in the metrics is maintained even in the DeiT training setting.